OpenStudy (anonymous):

Factor 4x2 – 13x + 10.

OpenStudy (anonymous):

Help @knock ?

OpenStudy (anonymous):

(x-2)(4x-5)=4x^2-5x-8x+10=4x^2-13x+10 (using Ruffini's rule)

OpenStudy (anonymous):

What's Ruffini's rule?

OpenStudy (jackellyn):

http://www.vitutor.com/alg/polynomials/ruffini.html Basically synthetic division.

OpenStudy (anonymous):

Nevermind I need help. Normal help.

OpenStudy (anonymous):

Use formula for factoring quadratic equation: ax2 + bx + c = a(x - x1)(x - x2) Where x1 and x2 are solutions of quadratic equation In our case we have: 4x2 - 13x + 10 = 4(x - 2)(x - 5/4) We can simplify this expression by multiply the second bracket by 4. After multiplying we have: 4x2 - 13x + 10 = ( x - 2 )( 4x - 5 )

OpenStudy (anonymous):

it's actually 4x^2 sorry.

OpenStudy (anonymous):

but still, the answer you get is (x-2)(4x-5)